OpenStudy (bethgrace96):

solve y=log (5x+3) for x

OpenStudy (solomonzelman):

\[ y=\log (5x+3)\]\[y~\log_{10}10 =\log (5x+3)\]\[\log_{10}10 ^y=\log (5x+3)\]\[10^y=5x+3\]\[10^y-3=5x\]\[\frac{10^y-3}{5}=x\]